Jump to content

Robe 150XT - Tilt problem


Recommended Posts

One of my Robe 150XT's has recently developed a tilt problem and after carrying out some diagnostic checks is proving to be a bit of a head scratcher.


The initial fault was that the tilt angle would randomly oscillate even though the DMX command was static. With the unit on the bench it is obvious that the stepper motor is not running smoothly and when sending different DMX levels to control the tilt angle it is a lottery as to which way the tilt will move and bears no relationship to the DMX level. Sometimes a tilt error is reported, other times not. Speaking to Robe UK they suspected that the problem was either due to a fault in the wiring harness to the stepper motor, opto sensor or the hall-effect switch or the sensor or switch itself. All of the wiring has now been checked out with both a meter and by isolating the connectors from the sockets and passing current through them just to ensure that the wiring has no fractures and is only continuous with a few strands of wire. These all check out fine.


I have swapped the opto sensor with the pan opto device and also verified the hall effect switch by monitoring the voltage on the respective pins when the magnet is close and moved away from the switch. Again all of these appear to be functioning correctly.


When adjusting the tilt via the internal test sequence in the unit I have noticed that depending upon the DMX value being sent the stepper motor has either normal or no control over the tilt movement. In other words if DMX level 4 is sent then the tilt motor has very little power and it is very easy to push the head to a different tilt angle with almost no force. However when a DMX level of 255 is being sent then the control from the tilt motor appears fairly normal and a reasonable amount of force is needed to push the head to a different angle as to be expected.


I am beginning to suspect that the problem is now related to the stepper motor driver ic and this has also been compounded by my observation that when I removed the pc board there was a some accumulation of dirt deposit on the back row of pins caused by the airflow driven through by the base cooling fan which may have contributed to the ic's demise. (In general the heads are fairly clean inside as I do try to keep them well maintained).


Before I start swapping the driver ic does anyone have any thoughts or similar experiences of this problem? It would also be useful to know if the driver ic is a stock item within RS for example as I cannot identify the device due to the integral heatsink.


Any pointers would be greatly appreciated. Thanks.



Link to comment
Share on other sites

The robe uses an L293D bridge for both pan & tilt, pwm for this comes from IC1. There is also an auxilliary pan & tilt power enable via fet T2, also controlled by IC1.


When fault finding a pan or tilt error, make it easy on yourself and turn the encoders off. This is done by PERS>FEED>OFF.


Then you can better see the 'real' fault rather than a tilt fault that may be partly masked by the encoder or a bad encoder making it look like a tilt fault.


The head should work perfectly with the encoders 'off' so long as you dont move the head with your hand after it's initialized.


I would change the wiring harness rather than mess about buzzing it through. Or at least thread a temporary set of tilt cables through to try it if you're that way inclined!


After the loom, I would suspect the motor for maybe weak magnetism. The pan & tilt drive ics are the same so you could always swap them I suppose.


The feel of the motor is bound to alter between doing small and large movements because the drive increases or decreases the motor current intelligently depending on where and how fast it calculates the motor needs to go and stop. It will also reduce the current to a holding level after a short period of non-movement.



Of course it may all work great with the encoders off and so that suggests an encoder error. Ditto re - wiring harness, dry joints etc which you say youve checked anyhow.


The tilt encoder signal is a 2-phase quadrature square wave present on pins 27 and 28 of IC2 and can be tested with an avo if you dont have a scope by operating the head slowly by hand with the tilt motor unplugged. (dont unplug it with power on).


Disable the lamp before operating it with the covers off, dont just unplug the lamp because the ignition voltage flashing over is just great for knocking out random functions on the PIC controllers.






Link to comment
Share on other sites

Thanks for the reply Kevin with some very helpful information.


The feedback option had already been switched off. With the encoders switched on it made the fault ten times worse and the head was in danger of shaking itself to bits as it was continually hunting for its correct position. The tilt motor has also been changed, in fact this was the first thing I did. The gobo stepper motor has the same part number and this was considered the easier option to swap rather than the pan motor.


I had also scoped the supplies to the two inputs of the stepper motor and this did seem to be giving me the sort of waveforms that I would have expected, so this was adding to the confusion.


Thanks for the part number of the driver ic. Given the minimal cost and ease of availability I think this has to be the next logical step to try. I'll report back after the driver has been changed.



Link to comment
Share on other sites


This topic is now archived and is closed to further replies.

  • Create New...

Important Information

We have placed cookies on your device to help make this website better. You can adjust your cookie settings, otherwise we'll assume you're okay to continue.